A circus tent is in the shape of a cone on the top of the cylinder whose height is 3 m and its diameter is 105 m, slant
height of the cone is 63 m. How much is the total area of the canvas used to make this tent?

Answer by mananth(16949)

total area of canvas = area of cylinder + area of cone
Area of cylinder = 2*pi*r*h
=2*pi*52.5*3
=989.6 sq.meters
..
area of cone = pi*r*s
=pi*52.5*63
=10390.81
total canvas needed = 989.6+10390.81
=11,380.41 sq.meters